OpenStudy (anonymous):

find an equation of the tangent to the curve at the given point. f(x)= 2sinx+3 at x=pi/6

OpenStudy (anonymous):

A tangent is a straight line. Therefore, the equation of a tangent can be written in the form y= mx +c where m=slope of the tangent c= y-intercept of the tangent

OpenStudy (anonymous):

to find the slope of the tangent at the point x=pi/6.. \[\frac{ dy }{ dx } (x=\frac{ \pi }{ 6 })\]

OpenStudy (anonymous):

do you get up to that point??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I plugged pi/6 on for all x

OpenStudy (anonymous):

yes....that is correct!

OpenStudy (anonymous):

then what do i do

OpenStudy (anonymous):

f(pi/6)=4

OpenStudy (anonymous):

that step is correct. you found the corresponding y-coordinate of the given x-coordinate (pi/6)

OpenStudy (anonymous):

y-4=m(x-pi/6) ??

OpenStudy (anonymous):

so now we can plug them in the equation y= mx +c to find the value of c. y=4 x=pi/6 m=sqrt of 3

OpenStudy (anonymous):

is that what you got for your value of the slope (m)??

OpenStudy (anonymous):

how do you get m=sqrt of 3??

OpenStudy (anonymous):

I differentiated the equation of the curve and replaced x with pi/6

OpenStudy (anonymous):

what did you get for the slope??

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

can you show me how you got m=sqrt of 3

OpenStudy (anonymous):

4=sqrt3(pi/6)+c ??

OpenStudy (anonymous):

you would have to know the basic differentials... \[y = \sin x \] \[\frac{ dy }{ dx } = \cos x\] \[y = \cos x \] \[\frac{ dy }{ dx } = - \sin x\] \[y = \tan x\] \[\frac{ dy }{ dx } = \sec ^{2}x\]

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are you ok with how i got the sqrt 3??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you can post that as a question because i use a shortcut method which i suppose would not be good for you....

OpenStudy (anonymous):

ok so you replaced sin x with cos x which is sqrt of 3/2

OpenStudy (anonymous):

yes....and you would have to know that if you differentiate a constant standing on its own you get zero...

OpenStudy (anonymous):

yes i understand

OpenStudy (anonymous):

ok....but i advice you would have to work on it...(on the topic differentiation)

OpenStudy (anonymous):

so what does c equal

OpenStudy (anonymous):

so have you found c??

OpenStudy (anonymous):

y-4=sqrt3(x-pi/6)

OpenStudy (anonymous):

if i have 4=sqrt3/6+c

OpenStudy (anonymous):

no..... \[4 = \sqrt{3} (\frac{ \pi }{ 6 }) + c\]

OpenStudy (anonymous):

so you subtract 4-sqrt3(pi/6)

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

are you supposed to leave your values in exact form or in decimals???

OpenStudy (anonymous):

exact form

OpenStudy (anonymous):

so what did you get as the value of c??

OpenStudy (anonymous):

3

OpenStudy (anonymous):

that would be a decimal

OpenStudy (anonymous):